# Context Parametrization with Compositional Adapters

### v2

@@ 6 Related Work
Analyses show ICL often exploits surface-level cues rather than learning task semantics (Min et al., 2022), and that transformers can internally implement simple learning algorithms during inference (Xie et al., 2022; Akyürek et al., 2023). In parallel, parameter-efficient fine-tuning (PEFT) replaces long prompts with compact trainable modules such as soft/prefix prompts (Li and Liang, 2021; Lester et al., 2021) or low-rank adapters (Hu et al., 2022). Recently, Hong et al. (2025) proposed mixtures of in-context learners to efficiently combine subsets of demonstrations. Like PEFT, our work keeps inference on the query only, but instead of concatenating demonstrations, we translate them into compositional adapters.
@@
Another line of research explores how to compose learned modules or entire models. In the adapter setting, AdapterFusion learns to fuse multiple task adapters non-destructively (Pfeiffer et al., 2021), while MAD-X composes language and task adapters for cross-lingual transfer (Pfeiffer et al., 2020). For full model weights, model soups average fine-tuned checkpoints to improve robustness (Wortsman et al., 2022), and task vectors use weight differences to add or subtract task behaviors (Ilharco et al., 2023). These approaches combine pre-trained modules or models after the fact. In contrast, we generate adapters on-the-fly from context and train them explicitly for additivity, so that parameter-space composition mirrors text concatenation.
@@ 7 Conclusion
We tackled the challenges of efficiency and instability that arise when adapting LLMs to tasks requiring multiple demonstrations or retrieved passages. To address these issues, we introduced CompAs, a meta-learning teacher–student framework that translates contexts into compositional adapters. CompAs encodes context into adapter parameters that can be algebraically combined, allowing processing of complex queries without input concatenation. In this way, we enable efficient handling of large context sets by generating adapters independently and composing them in parameter space, reducing inference cost, mitigating long-context degradation, and circumventing context window limitations. A reconstruction objective promotes safety and security, ensuring that the input context is decodable from the adapter parameters. CompAs consistently outperforms ICL and prior generator-based approaches on multi-choice and extractive question answering. Taken together, these results establish composable context parametrization as a scalable approach for adapting LLMs.
Reproducibility Statement
Along with the description of our method in Section 3 and the experimental setup in Appendix B, we also provide comprehensive details to ensure reproducibility. All datasets used (MMLU, ARC-Challenge, HotpotQA) are publicly available and described in Section 4.1. Model configurations, optimization settings, and training schedules are fully specified in Appendix B, along with the specified computing infrastructure. Theoretical results, including the proof of Theorem 1 and its corollary, are presented in Appendix A. Prompt templates for all benchmarks are provided in Appendix D. In addition, we include code in the supplementary material and will publish the repository publicly after the review process. Collectively, these resources allow independent reproduction and verification of our experiments and analyses.
@@ Appendix A Compositionality Bound
A.1 Proof of Theorem 1
Let $\bm{\phi}_{1}=G(\mathbf{c}_{1})$, $\bm{\phi}_{2}=G(\mathbf{c}_{2})$, and $\bm{\phi}_{12}=G([\mathbf{c}_{1};\mathbf{c}_{2}])$. We aim to bound
$\big\|\,\mathbf{f}_{S}^{\bm{\phi}_{1}+\bm{\phi}_{2}}(\mathbf{q})-\mathbf{f}_{T}([\mathbf{c}_{1};\mathbf{c}_{2};\mathbf{q}])\,\big\|_{2}.$
@@ A.1 Proof of Theorem 1
$a=\mathbf{f}_{S}^{\bm{\phi}_{1}+\bm{\phi}_{2}}(\mathbf{q}),\quad b=\mathbf{f}_{S}^{\bm{\phi}_{12}}(\mathbf{q}),\quad c=\mathbf{f}_{T}([\mathbf{c}_{1};\mathbf{c}_{2};\mathbf{q}]),$
@@
$\|\mathbf{f}_{S}^{\bm{\phi}_{1}+\bm{\phi}_{2}}(\mathbf{q})-\mathbf{f}_{T}([\mathbf{c}_{1};\mathbf{c}_{2};\mathbf{q}])\|_{2}$
$\qquad\leq\underbrace{\|\mathbf{f}_{S}^{\bm{\phi}_{1}+\bm{\phi}_{2}}(\mathbf{q})-\mathbf{f}_{S}^{\bm{\phi}_{12}}(\mathbf{q})\|_{2}}_{\text{Term (I)}}\;+\;\underbrace{\|\mathbf{f}_{S}^{\bm{\phi}_{12}}(\mathbf{q})-\mathbf{f}_{T}([\mathbf{c}_{1};\mathbf{c}_{2};\mathbf{q}])\|_{2}}_{\text{Term (II)}}.$ (6)
@@
$\text{Term (I)}\;\leq\;L\,\|\bm{\phi}_{1}+\bm{\phi}_{2}-\bm{\phi}_{12}\|_{2}\;=\;L\,\eta.$ (7)
@@
$\|\mathbf{f}_{S}^{\bm{\phi}_{1}+\bm{\phi}_{2}}(\mathbf{q})-\mathbf{f}_{T}([\mathbf{c}_{1};\mathbf{c}_{2};\mathbf{q}])\|_{2}\;\leq\;L\eta+\epsilon([\mathbf{c}_{1};\mathbf{c}_{2}]),$
which proves Theorem 1. ∎
Corollary 1 (Extension to $k$ contexts).
By induction, Theorem 1 extends to any sequence of contexts $\mathbf{c}_{1},\ldots,\mathbf{c}_{k}$ with corresponding adapters $\bm{\phi}_{i}=G(\mathbf{c}_{i})$:
$\|\mathbf{f}_{S}^{\sum_{i=1}^{k}\bm{\phi}_{i}}(\mathbf{q})-\mathbf{f}_{T}([\mathbf{c}_{1};\ldots;\mathbf{c}_{k};\mathbf{q}])\|_{2}\;\leq\;(k-1)L\eta\;+\;\epsilon([\mathbf{c}_{1};\ldots;\mathbf{c}_{k}]).$
@@ A.2 Geometric interpretation
Adapters $\bm{\phi}=G(\mathbf{c})$ can be viewed as vectors in the parameter space $\Phi$. In the ideal case, $G$ embeds contexts into a flat subspace where concatenation in input space corresponds exactly to vector addition in parameter space. The deviation terms then acquire a geometric meaning: $\eta$ measures the generator’s departure from additivity, and $\epsilon([\mathbf{c}_{1};\mathbf{c}_{2}])$ captures the student’s misalignment with the teacher on concatenated contexts. The Lipschitz constant $L$ governs how deviations in adapter space propagate into the model’s output space. Compositionality can therefore be understood as flattening the contextual geometry into a nearly linear embedding, with small $\eta$ and $\epsilon$ ensuring that adapter addition faithfully mirrors context concatenation.
@@ B.4 Hyperparameters
We insert LoRA adapters into all attention projections ($q$, $k$, $v$, and $o$) as well as into the MLP down- and up-projection layers. For the generator, we use a uniform rank of $32$ across all modules, while for the student we use rank $16$. Following standard practice, each LoRA module applies a scaling factor $\alpha$, such that the effective weight update is $\frac{\alpha}{r}AB$. We set $\alpha=32$ for generator adapters and $\alpha=16$ for student adapters, ensuring balanced contribution of low-rank updates relative to their respective ranks.
@@ B.5 Computing Infrastructure
We conducted our experiments on a mix of local and cluster resources. Local training was performed on a workstation with an AMD Ryzen Threadripper 3970X 32-Core CPU, 256GB RAM, and $2\times$ NVIDIA GeForce RTX 3090 GPUs (24GB each). Larger-scale runs were executed on a compute cluster equipped with $2\times$ NVIDIA A100 GPUs (40GB each).
@@ C.3 Positional Encoding and Order Sensitivity
By design, our generator enforces commutativity, i.e., $G([c_{1};c_{2}])\approx G(c_{1})+G(c_{2})=G(c_{2})+G(c_{1})$, which is desirable for tasks where order invariance is beneficial and directly reduces $\eta$. The trade-off is that strict commutativity removes the ability to encode order.
As a proof of concept, we conduct an experiment on HotpotQA using pairs of supporting contexts. We augment the generator with slot embeddings, i.e., for slot $i$ we add a learnable vector $s_{i}$ to the corresponding adapter delta:
$\tilde{\phi}_{i}=G(c_{i})+s_{i},\qquad\Phi=\sum_{i}\tilde{\phi}_{i}.$ (9)
This introduces positional variance while preserving linear composition, thereby allowing the model to distinguish $c_{1};c_{2}]$ from $c_{2};c_{1}]$.
@@
LLaMA-3.1 8B LLaMA-3.2 3B Qwen-2.5 7B No-PE (commutative linear) $62.5_{2.0}$ $50.1_{2.2}$ $60.4_{2.1}$ Slot-PE $87.6_{0.9}$ $71.1_{2.6}$ $85.0_{1.4}$
@@ C.4 Compositional Scaling
Total demos Method Accuracy $\bm{\eta}$ $\downarrow$
@@ C.6 Efficiency
Table 13: Absolute compute comparison with LLaMA 8B. We report absolute inference-time compute per query and training compute in TFLOPs on the ARC-Challenge dataset using $16$ demonstrations in total with $4\times 4$ composition for CompAs. WILDA trains a new adapter for each demonstration configuration, whereas GenAda and CompAs train a generator once and amortize across different contexts.
Method Inference TFLOPs / query Training TFLOPs Regime ICL (prompting) $4.0$ No training WILDA $0.7$ $1.4\times 10^{5}$ Per demo set GenAda $0.9$ $3.1\times 10^{7}$ One-time CompAs $0.8$ $2.4\times 10^{6}$ One-time
@@ Appendix E Limitations
LLM Usage Disclosure
In preparing this manuscript, we used LLMs to polish the writing at the sentence level. This assistance was limited to stylistic improvements such as grammar, clarity, and conciseness.

### v3

@@ 6 Related Work
Analyses of ICL suggest that models often exploit surface-level cues rather than learning task semantics (Min et al., 2022), and that transformers can internally implement simple learning algorithms during inference (Xie et al., 2022; Akyürek et al., 2023). In parallel, parameter-efficient fine-tuning (PEFT) replaces long prompts with compact trainable modules such as soft/prefix prompts (Li and Liang, 2021; Lester et al., 2021) or low-rank adapters (Hu et al., 2022). Recently, Hong et al. (2025) proposed mixtures of in-context learners to efficiently combine subsets of demonstrations. Like PEFT, our work keeps inference on the query only, but instead of concatenating demonstrations, we translate them into compositional adapters.
@@
Another line of research explores how to compose learned modules or entire models. In the adapter setting, AdapterFusion learns to fuse multiple task adapters non-destructively (Pfeiffer et al., 2021), while MAD-X composes language and task adapters for cross-lingual transfer (Pfeiffer et al., 2020). For full model weights, model soups average fine-tuned checkpoints to improve robustness (Wortsman et al., 2022), and task vectors use weight differences to add or subtract task behaviors (Ilharco et al., 2023). These approaches combine pre-trained modules or models after the fact. In contrast, we generate adapters on-the-fly from context and train them explicitly for additivity, such that their composition in parameter space is functionally equivalent to text concatenation.
@@ 7 Conclusion
We tackled the challenges of efficiency and instability that arise when adapting LLMs to tasks requiring multiple demonstrations or retrieved passages. To address these issues, we introduced CompAs, a meta-learning teacher–student framework that translates contexts into compositional adapters. CompAs encodes context into adapter parameters that can be algebraically combined, allowing processing of complex queries without input concatenation. In this way, we enable efficient handling of large context sets by generating adapters independently and composing them in parameter space, reducing inference cost, mitigating long-context degradation, and circumventing context window limitations. A reconstruction objective promotes safety and security, ensuring that the input context is decodable from the adapter parameters. CompAs consistently outperforms ICL and prior generator-based approaches on multi-choice and extractive question answering. Taken together, these results establish compositional context parametrization as a scalable approach for adapting LLMs.
Impact Statement
This work proposes encoding contextual information into compositional adapter parameters rather than using long input prompts. Such modular parameterization can reduce inference costs, improve stability in long-context settings, and enable more efficient adaptation of large language models. By structuring context as components in parameter space, the approach may facilitate modular updates and systematic analysis compared to prompt-based adaptation. However, encoding context into parameters introduces privacy and security considerations. If sensitive information is used during training, it may become recoverable from learned adapters. Practical deployment should therefore incorporate safeguards such as controlled access to adapter parameters, data minimization, and privacy-preserving training methods.
@@ Appendix A Compositionality Bound
A.1 Proof of Theorem 2.1
Let $\boldsymbol{\phi}_{1}=G(\mathbf{c}_{1})$, $\boldsymbol{\phi}_{2}=G(\mathbf{c}_{2})$, and $\boldsymbol{\phi}_{12}=G([\mathbf{c}_{1};\mathbf{c}_{2}])$. We aim to bound
$\big\|\,\mathbf{f}_{S}^{\boldsymbol{\phi}_{1}+\boldsymbol{\phi}_{2}}(\mathbf{q})-\mathbf{f}_{T}([\mathbf{c}_{1};\mathbf{c}_{2};\mathbf{q}])\,\big\|_{2}.$
@@
$a=\mathbf{f}_{S}^{\boldsymbol{\phi}_{1}+\boldsymbol{\phi}_{2}}(\mathbf{q}),\quad b=\mathbf{f}_{S}^{\boldsymbol{\phi}_{12}}(\mathbf{q}),\quad c=\mathbf{f}_{T}([\mathbf{c}_{1};\mathbf{c}_{2};\mathbf{q}]),$
@@
$\|\mathbf{f}_{S}^{\boldsymbol{\phi}_{1}+\boldsymbol{\phi}_{2}}(\mathbf{q})-\mathbf{f}_{T}([\mathbf{c}_{1};\mathbf{c}_{2};\mathbf{q}])\|_{2}$
$\qquad\leq\underbrace{\|\mathbf{f}_{S}^{\boldsymbol{\phi}_{1}+\boldsymbol{\phi}_{2}}(\mathbf{q})-\mathbf{f}_{S}^{\boldsymbol{\phi}_{12}}(\mathbf{q})\|_{2}}_{\text{Term (I)}}\;+\;\underbrace{\|\mathbf{f}_{S}^{\boldsymbol{\phi}_{12}}(\mathbf{q})-\mathbf{f}_{T}([\mathbf{c}_{1};\mathbf{c}_{2};\mathbf{q}])\|_{2}}_{\text{Term (II)}}.$ (6)
@@
$\text{Term (I)}\;\leq\;L\,\|\boldsymbol{\phi}_{1}+\boldsymbol{\phi}_{2}-\boldsymbol{\phi}_{12}\|_{2}\;=\;L\,\eta.$ (7)
@@
$\|\mathbf{f}_{S}^{\boldsymbol{\phi}_{1}+\boldsymbol{\phi}_{2}}(\mathbf{q})-\mathbf{f}_{T}([\mathbf{c}_{1};\mathbf{c}_{2};\mathbf{q}])\|_{2}\;\leq\;L\eta+\epsilon([\mathbf{c}_{1};\mathbf{c}_{2}]),$
which proves Theorem 2.1. ∎
Corollary A.1 (Extension to $k$ contexts).
By induction, Theorem 2.1 extends to any sequence of contexts $\mathbf{c}_{1},\ldots,\mathbf{c}_{k}$ with corresponding adapters $\boldsymbol{\phi}_{i}=G(\mathbf{c}_{i})$:
$\|\mathbf{f}_{S}^{\sum_{i=1}^{k}\boldsymbol{\phi}_{i}}(\mathbf{q})-\mathbf{f}_{T}([\mathbf{c}_{1};\ldots;\mathbf{c}_{k};\mathbf{q}])\|_{2}\;\leq\;(k-1)L\eta\;+\;\epsilon([\mathbf{c}_{1};\ldots;\mathbf{c}_{k}]).$
@@ A.2 Geometric interpretation
Adapters $\boldsymbol{\phi}=G(\mathbf{c})$ can be viewed as vectors in the parameter space $\Phi$. In the ideal case, $G$ embeds contexts into a flat subspace where concatenation in input space corresponds exactly to vector addition in parameter space. The deviation terms then acquire a geometric meaning: $\eta$ measures the generator’s departure from additivity, and $\epsilon([\mathbf{c}_{1};\mathbf{c}_{2}])$ captures the student’s misalignment with the teacher on concatenated contexts. The Lipschitz constant $L$ governs how deviations in adapter space propagate into the model’s output space. Compositionality can therefore be understood as flattening the contextual geometry into a nearly linear embedding, with small $\eta$ and $\epsilon$ ensuring that adapter addition faithfully mirrors context concatenation.
@@ B.4 Hyperparameters
We insert LoRA adapters into all attention projections ($q$, $k$, $v$, and $o$) as well as into the MLP down- and up-projection layers. For the generator, we use a uniform rank of $32$ across all modules, while for the student, we use rank of $16$. Following standard practice, each LoRA module applies a scaling factor $\alpha$, such that the effective weight update is $\frac{\alpha}{r}AB$. We set $\alpha=32$ for generator adapters and $\alpha=16$ for student adapters, ensuring balanced contribution of low-rank updates relative to their respective ranks.
@@ B.5 Computing Infrastructure
We conducted our experiments on a mix of local and cluster resources. Local training was performed on a workstation with an AMD Ryzen Threadripper 3970X 32-Core CPU, 256GB RAM, and NVIDIA GeForce RTX 3090 GPUs. Larger-scale runs were executed on a compute cluster equipped with NVIDIA A100 GPUs.
@@ C.3 Positional Encoding and Order Sensitivity
By design, our generator enforces commutativity, i.e., $G([\mathbf{c}_{1};\mathbf{c}_{2}])\approx G(\mathbf{c}_{1})+G(\mathbf{c}_{2})=G(\mathbf{c}_{2})+G(\mathbf{c}_{1})$, which is desirable for tasks where order invariance is beneficial and directly reduces $\eta$. The trade-off is that strict commutativity removes the ability to encode order.
As a proof of concept, we conduct an experiment on HotpotQA using pairs of supporting contexts. We augment the generator with slot embeddings, i.e., for slot $i$ we add a learnable vector $\mathbf{s}_{i}$ to the corresponding adapter delta:
$\tilde{\boldsymbol{\phi}}_{i}=G(\mathbf{c}_{i})+\mathbf{s}_{i},\qquad\boldsymbol{\phi}=\sum_{i}\tilde{\boldsymbol{\phi}}_{i}.$ (9)
This introduces positional variance while preserving linear composition, thereby allowing the model to distinguish $\mathbf{c}_{1};\mathbf{c}_{2}]$ from $\mathbf{c}_{2};\mathbf{c}_{1}]$.
@@
LLaMA-3.1 8B LLaMA-3.2 3B Qwen-2.5 7B
No-PE (commutative linear) $62.5_{2.0}$ $50.1_{2.2}$ $60.4_{2.1}$
Slot-PE $87.6_{0.9}$ $71.1_{2.6}$ $85.0_{1.4}$
@@ C.4 Compositional Scaling
Total demos Method Accuracy $\boldsymbol{\eta}$ $\downarrow$
@@ C.6 Efficiency
Table 13: Absolute compute comparison with LLaMA 8B. We report absolute inference-time compute per query and training compute in TFLOPs on the ARC-Challenge dataset using $16$ demonstrations in total with $4\times 4$ composition for CompAs. CompAs trains a new adapter for each demonstration configuration, whereas GenAda and CompAs train a generator once and amortize across different contexts.
Method Inference TFLOPs / query Training TFLOPs Regime
ICL (prompting) $4.0$ No training
wilda $0.7$ $1.4\times 10^{5}$ Per demo set
GenAda $0.9$ $3.1\times 10^{7}$ One-time
CompAs $0.8$ $2.4\times 10^{6}$ One-time
@@ Appendix E Limitations
Appendix F Reproducibility Statement
Along with the description of our method in Section 3 and the experimental setup in Appendix B, we provide comprehensive details to ensure reproducibility. All datasets used (MMLU, ARC-Challenge, HotpotQA) are publicly available and described in Section 4.1. Model configurations, optimization settings, and training schedules are fully specified in Appendix B, along with the specified computing infrastructure. Theoretical results, including the proof of Theorem 2.1 and its corollary, are presented in Appendix A. Prompt templates for all benchmarks are provided in Appendix D. In addition, we include code in the supplementary material and will publish the repository publicly after the review process.